Grade 12th passPhysical Chemistry

x mmol of KIO3 reacts completely with y mmol of KI to give I2 quantitatively. If z mmol of hypo are required for complete titration against this I2 then which relation is not correct ?
a. z = 6x
b. 6y = 5z
c. x = 5y
d. x + y = 2z
please please give detailed answer.....

Yasir Rather
7 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
Here, KIO3 + 5KI + 3H2SO4 → 3I2 + 3H2O + 3K2SO4 And 6Na2S2O3 + 3I2 → 3Na2S4O6 + 6NaI From the above two chemical equation, we get : x = 1, y = 5 and z = 6 Now, KIO3 + 6Na2S2O3 + 3H2SO4 → 3NaS4O6 + KI + 3H2O + 3Na2SO4 We have , Coa (Concentration of oxidising agent) x Voa (volume of oxidising agent)/ 1 = Cra x Vra (Reducing agent) / 6 = > x/1 = z/6 [v = 1 (assumed)] => z = 6x Similarly, when we will have KI as the oxidising agent and Na2S2O3 as the reducing agent : y/5 = z/6 => 6y = 5z Now, x + y = (1 + 5) = 6 = z so, option (c) is not correct.